### v1

# Acceleration of positive muons by a radio-frequency cavity

Abstract
Acceleration of positive muons from thermal energy to 100 keV has been demonstrated. Thermal muons were generated by resonant multi-photon ionization of muonium atoms emitted from a sheet of laser-ablated aerogel. The thermal muons were first electrostatically accelerated to 5.7 keV, followed by further acceleration to 100 keV using a radio-frequency quadrupole. The transverse normalized emittance of the accelerated muons in the horizontal and vertical planes were mmmrad and mmmrad, respectively. The measured emittance values demonstrated phase space reduction by a factor of (horizontal) and (vertical) allowing good acceleration efficiency. These results pave the way to realize the first-ever muon accelerator for a variety of applications in particle physics, material science, and other fields.
An accelerated positive muon beam enables a variety of next-generation muon experiments. One such application is a new experiment to measure the muon anomalous magnetic moment and search for the muon electric dipole moment [1]. A transmission muon microscope has also been proposed for high-resolution imaging of thick materials, which is not achievable with conventional transmission electron microscopes [2]. A or collider is another promising application of the beam for the energy frontier in the future [3].
A high-intensity muon beam with more than 108 muons per second can be produced from the decay of pions generated in a pion production target irradiated by a MW-class proton beam. However, as a tertiary beam, the muon beam initially has a large phase-space volume. Cooling the muon beam to increase its phase-space density is necessary to realize an efficient acceleration of muons. On the other hand, applying traditional cooling techniques, such as synchrotron radiation cooling, laser cooling [4], stochastic cooling [5], and electron cooling [6], to muons is not feasible due to the short lifetime of muons.
The cooling techniques dedicated to muons have been intensively developed. The ionization cooling of both positive and negative muons was proposed for a muon collider [7]. The transverse and longitudinal momentum of the muon beam is lost inside an absorber through ionization, and the longitudinal momentum is restored using radio-frequency cavities. cooling positive muons inside cryogenic helium gas combined with a complex electromagnetic field was also proposed [8]. This technique aims to improve the phase space density of a muon beam by a factor of 1010 with the efficiency of 10-3. These cooling techniques have been demonstrated recently [9, 10, 11]. However, the subsequent acceleration of muons after the cooling has not been reported so far.
Another technique of muon cooling and acceleration has been under development at the muon science facility (MUSE) [12] of the Japan Proton Accelerator Research Complex (J-PARC). This technique cools positive muons down to the thermal energy and accelerates them using radio-frequency (rf) cavities. The thermal energy muon produced from this technique is called ultra-slow muon. The ultra-slow muons are generated through resonant multi-photon ionization of thermal muonium atoms () emitted from a muonium production target irradiated by a conventional muon beam [13, 14]. Normalized transverse root-mean-square (rms) emittance as low as 0.3 mmmrad is possible with this technique at J-PARC [1]. The ultra-slow muons are extracted and transported to linear accelerators by an electrostatic field. Due to the simple extraction of the cooled muons, this technique is a promising approach to accelerate the cooled muons.
In this letter, the first rf-acceleration of ultra-slow muons to 100 keV is reported. The emittance of the accelerated muon beam is also evaluated using the quadrupole scan method.
The experiment reported here was conducted in the S2 area, one of the experimental areas of the surface muon beamline (S-line) at MUSE, during March and April of 2024. Figure 1 shows the experimental setup for the muon acceleration, consisting of a silica aerogel target for muonium production, an ionization laser, an electrostatic lens (Soa lens), electrostatic deflectors, a radio-frequency quadrupole (RFQ), and a beam diagnostic line.
The MUSE facility provides a pulsed surface muon beam produced by decay near the surface of the production target with a repetition rate of 25 Hz. In this experiment, the proton beam power was 940 kW, and the intensity was muons per pulse (/pulse). The profile of the surface muon beam was measured by a muon beam monitor with a gated image intensifier [15]. The horizontal and vertical rms beam sizes were 24 and 13 mm, respectively. The mean and the rms value of the momentum were 28 MeV/ and 0.46 MeV/, respectively, based on the simulation of the S-line beam transport using the G4BEAMLINE software [16]. The normalized transverse rms emittance () in the horizontal and vertical plane were mmmrad and mmmrad, respectively, based on the simulation.
The surface muons were decelerated by a 420 m-thick aluminum foil and then incident on a silica aerogel target. The target had a diameter of 78 mm, a thickness of 8 mm, and a density of 23.6 mg/cm3. Both sides of the target were ablated by a short pulse laser, creating 250 m diameter holes at 300 m intervals [17, 18]. A mesh electrode was placed to the downstream surface of the target for the extraction of the ionized muons. The stopping efficiency of the surface muon inside the target was 38% evaluated by a simulation based on Geant4 [19]. The conversion efficiency from muon to muonium was 52% [20]. The total muonium emission efficiency from the target was 13% from a simulation based on a three-dimensional random walk model of muonium diffusion inside the target [21]. Of the emitted muonium atoms, 19% were lost at the mesh-electrode.
A narrow-band pulsed light source operated at a wavelength of 244 nm, originally developed for the muonium 1S-2S spectroscopy experiment [22], was introduced to first excite the muonium atoms to the 2S state and then ionize them from that state. The light source was based on an injection-seeded titanium-sapphire laser operated at a wavelength of 976 nm. The titanium-sapphire crystals were pumped by an all-solid-state light source operated at a wavelength of 532 nm. The energy of the 244 nm pulse was 0.5–1.3 mJ, depending on the alignment and damage of the optics. The beam diameter was 1 mm. The pulse duration was 57 ns. The pulse traveled horizontally and parallel to the aerogel surface at a 2 mm distance and was reflected by a mirror for Doppler-free 1S-2S transition. The spatial distribution of the ultra-slow muon was highly anisotropic as the beam size along the vertical axis was determined by the radius of the laser, while the size along the horizontal axis was determined by the size of the surface muon beam. The ionization efficiency was defined as the ratio of the ionized muonium atoms to the total emitted muonium atoms, including the muon decay loss after the emission. Based on simulations, the efficiency was – , depending on the pulse energy.
The ultra-slow muons were accelerated to 5.7 keV and focused at the entrance of the RFQ by an electrostatic lens system, a Soa lens [23]. The extraction energy and focal length were determined to match the acceptance of the following acceleration using the musrSim [24], which is based on Geant4. Three pairs of Helmholtz coils were introduced to compensate for the ambient magnetic field around the muonium target in three directions. Four sets of electrostatic deflectors were used to correct the muon trajectory deflected by the ambient magnetic field. 20% of the muons were lost by the muon decay during the transport based on a simulation.
The intensity of the 5.7 keV muon before the rf-acceleration was /pulse. For this measurement, a slow muon beam line (SMBL) [14] was installed after the Soa lens instead of the deflector and downstream components. The laser frequency was scanned and tuned to excite the 1S-2S transition of muonium in the state. The laser irradiation timing was also scanned to maximize the intensity of the muon signal and fixed to 1.25 s after the arrival time of the surface muons to the target.
A prototype RFQ of the J-PARC linac [25] was used for the acceleration of the ultra-slow muon. The RFQ was designed to accelerate negative hydrogen ions up to 0.8 MeV. This RFQ also accelerated negative muonium atoms up to 89 keV [26, 27]. The forward rf frequency, peak power, and pulse width were set to 324.03 MHz, 2.6 kW, and 40 s, respectively. For the beam dynamics simulation of the RFQ, PARMTEQM [28] was used. The transmittance of the RFQ was 81% based on a simulation, mostly due to decay losses.
The beam properties of accelerated muons were measured using a beam diagnostics line after the RFQ [29]. The beam was transferred using two quadrupole magnets (QM1 and QM2) with field gradients of , where is the current supplied to each QM. The charge and momentum of the particle were selected using a horizontal bending magnet (BM). The beam transportation was simulated using General Particle Tracer [30]. The beam loss due to the muon decay was 3%.
A single-anode microchannel plate (MCP) located at the downstream end of the beam diagnostics line was used to measure the arrival time of the particles. The effective area of the MCP corresponded to a circle 42 mm in diameter, with an aperture ratio of 60%. The output signal of the MCP was digitized with a 250 MS/s waveform digitizer. For the profile measurement of muons, an MCP-based beam profile monitor (BPM) was used [31]. The effective area of the BPM was 40 mm in diameter. The exposure time of the profile monitor was 500 ns for each muon pulse.
Figure 2 shows the time distributions of the MCP pulses where the time origin is the arrival of the surface muon at the muonium target. A clear peak was observed at 2280 ns when both the rf power was turned on, and the laser frequency was tuned to the resonance of the 1S-2S transition, corresponding to the accelerated muons. The time of flight (TOF), defined as the time difference between the laser irradiation and the muon arrival at the MCP, was ns, where the first uncertainty was statistical and the second was systematic, mainly coming from the uncertainty of the laser irradiation time. The TOF agrees with the simulated value of ns, where the uncertainty came from the muon transportation inside the Soa lens. The muon intensity was /pulse during this measurement, with 10% – 40% smaller pulse energy of the laser compared to the pulse energy during the measurement of the 5.7 keV muon intensity.
The measurement of for the accelerated muon beam was conducted using the quadrupole scan method [32]. The rms beam sizes of the muon beam after the BM were measured for different quadrupole strengths of QM2. Figure 3 shows an example of the measured muon beam profile when A and A, measured during the quadrupole scan for in the vertical plane. The rms of the beam profile along the horizontal (x) and vertical (y) axis were calculated from the signal distribution. The signal-to-noise ratio, approximately 50 based on the measurements taken under the laser off-resonance condition, was accounted for in the rms calculation. In the case of the profile shown in Fig. 3, the rms beam size along the y-axis was mm, where the first uncertainty was statistical and second systematic. The systematic uncertainty, originating from the uncertainty in the number of background events, was evaluated using a Monte Carlo simulation.
Figure 4 shows the results of the quadruple scan. The normalized rms emittance, , was estimated by fitting them using the transfer matrix, which included QM1, QM2, BM, and the drift space. Data points corresponding to beam sizes larger than 5 mm were excluded from the fit, as larger beam sizes could introduce bias in the rms calculation due to detector size limitations. The momentum spread of the muon beam, evaluated to be from PARMTEQM, was included in the fitting function to account for the dispersion. The uncertainty of the momentum spread originated from the rf-power evaluation.
Systematic uncertainties on are shown in Table 1. The uncertainty from the momentum spread was assessed by varying its value in the fitting function. The uncertainty from beam loss in the diagnostic line was evaluated through a simulation.
Finally, the measured of the accelerated muon beam in the horizontal and vertical planes were mmmrad and mmmrad, respectively. The simulated in the horizontal and vertical plane were mmmrad and mmmrad, respectively. The origin of the discrepancies between measured and simulated emittance was inferred as an unaccounted-for mismatch factor in the simulation.
| Source | Contribution [ mmmrad] | |
| Horizontal | Vertical | |
| Rms beam size | ||
| Momentum spread | – | |
| Beam loss | ||
| Total | ||
In summary, muons were cooled and accelerated by an rf accelerator for the first time. The ultra-slow muons were generated by the resonant multi-photon ionization of muonium atoms emitted from a laser-ablated aerogel target using an all-solid-state laser at 244 nm, and accelerated to 100 keV. Compared to the initial normalized transverse rms emittance of the surface muon beam, the amounts of the phase-space reductions were by a factor of (horizontal) and (vertical), showing the realization of a low emittance muon beam accelerated by an rf-cavity.
The results presented in this paper show the phase space reduction of a muon beam by more than two orders of magnitude. This enables broad applications of an accelerated positive muon beam from particle physics to various other scientific fields.
Acknowledgments
The authors would like to thank the J-PARC muon section staff for their support in the conduct of the experiment at J-PARC MUSE. This work was supported in part by JSPS Kakenhi Grant Numbers JP18H05226, JP19H05606, JP20H05625, JP21K13944, JP21J01132, JP22KJ1594, JP22K21350, JP22H00141, JP24H00023, JP24K03211, MEXT Q-LEAP JPMXS0118069021, and JST-Mirai Program JPMJMI17A1. The experiment was performed at the Materials and Life Science Experimental Facility of the J-PARC under a user program 2011MS06.